# Continual Unsupervised Out-of-Distribution Detection

### v1

@@ Abstract
Deep learning models excel when the data distribution during training aligns with testing data. Yet, their performance diminishes when faced with out-of-distribution (OOD) samples, leading to great interest in the field of OOD detection. Current approaches typically assume that OOD samples originate from an unconcentrated distribution complementary to the training distribution. While this assumption is appropriate in the traditional unsupervised OOD (U-OOD) setting, it proves inadequate when considering the place of deployment of the underlying deep learning model. To better reflect this real-world scenario, we introduce the novel setting of continual U-OOD detection. To tackle this new setting, we propose a method that starts from a U-OOD detector, which is agnostic to the OOD distribution, and slowly updates during deployment to account for the actual OOD distribution. Our method uses a new U-OOD scoring function that combines the Mahalanobis distance with a nearest-neighbor approach. Furthermore, we design a confidence-scaled few-shot OOD detector that outperforms previous methods. We show our method greatly improves upon strong baselines from related fields.
@@
Out-of-distribution detection Continual learning
@@ 1 Introduction
Deep learning (DL) models fundamentally rely on the premise that the data distribution during training is congruent with that of the testing data. This assumption, when valid, enables these models to excel across a multitude of domains. However, this assumption often fails in real-world situations beyond the confines of controlled academic settings, primarily due to the emergence of unexpected samples, such as those stemming from issues with image acquisition. For these out-of-distribution (OOD) samples, the performance of DL models cannot be guaranteed, nor can they be readily identified. As a result, the dependability of DL models in applications where safety is paramount remains uncertain.
These shortcomings have sparked great interest in the field of OOD detection [17, 52]. OOD detectors are tasked with identifying and removing any sample that the underlying DL model can not reliably process, thereby increasing overall safety. Within the field, a natural distinction can be made between supervised and unsupervised OOD detection. Supervised OOD detectors use either a labeled training dataset or explicit OOD samples to model the training distribution. In contrast, unsupervised OOD (U-OOD) detection assumes neither access to training labels nor OOD samples, leading to a more generally applicable albeit more challenging setting.
Figure 1: The continual unsupervised out-of-distribution detection setting. In the first step, a U-OOD detector is trained using only unlabeled in-distribution data. Subsequently, during deployment, the OOD detector encounters real samples from both the in-distribution and out-distribution. Without knowing their categorization, these samples are integrated into the continual U-OOD detector, enhancing its ability to capture the actual OOD distribution accurately.
The same assumption that underpins closely related fields such as anomaly detection is used in current U-OOD works: normal samples are concentrated, whereas OOD samples are not [46, 48]. This reduces the problem of U-OOD to finding the level sets of the training data distribution and setting a threshold to define OOD samples. However, this assumption fails to consider the underlying motivation of OOD detection, that is, improving the reliability of downstream models. As U-OOD detectors cannot rely on training labels for context [1], we argue they must infer this from their deployment setting. In turn, the deployment context fully determines a concentrated OOD distribution.
To illustrate this, consider a U-OOD detector trained on a dataset of chest X-rays and intended for deployment in conjunction with a certain X-ray scanner. Within this context, the likelihood of encountering faulty images from scanner mishandling is magnitudes higher than encountering an image from a telescope or an image of the ocean. Furthermore, the exact composition of the OOD distribution depends entirely on external factors such as the scanner type and user proficiency. As a result, U-OOD works that synthesize anomalies (e.g., [45]) to simulate the OOD distribution are unable to capture this. Instead, an approach where an initial U-OOD detector, agnostic to the OOD distribution, is gradually updated during deployment to consider the actual OOD distribution is needed in practice (see Fig. 1).
@@
Yet, current U-OOD research has mainly neglected this important consideration. Most closely related to this topic are the fields of OOD detection with in-the-wild data [8, 23] and OOD test-time adaptation [10, 54]. The former assumes access to a large set of unlabeled data consisting of both in-distribution (ID) and OOD samples, whereas the latter uses the entire unlabeled test set to update a model. In both cases, hundreds of OOD samples are used to update their respective models. Consequently, this is often unrealistic, considering that OOD samples are scarce in real-world settings.
@@
Instead, we focus here on improving detection with only a few OOD samples and tackle the setting where this process is iterative: detect OOD samples, update the OOD detector, and subsequently repeat the procedure with a refined model several times. Consequently, evaluating U-OOD detectors should extend beyond a single instance and consider their effectiveness over time. To do this, we propose the setting of continual U-OOD detection, where detectors are evaluated over multiple timesteps. In the first timestep, a standard U-OOD detector is trained using unlabeled in-distribution data only. For every subsequent timestep, the model is deployed to assess unlabeled test samples, including a limited number of out-of-distribution samples, which need to be incorporated into the model. We believe this setting most closely mimics that of the real world. As such, the main contributions of the present work are:
@@
    We introduce the problem of continual U-OOD detection, which closely matches the reality of model deployment. By doing so, we introduce new metrics and benchmarks to evaluate methods in this setting.
@@
    We propose a confidence-scaled few-shot OOD detector, which uses a U-OOD scoring function combining the Mahalanobis distance with a nearest-neighbor approach.
@@
    We propose a novel method for continual U-OOD detection that outperforms strong baselines from multiple related fields.
2 Related Work
The vast majority of OOD works fall into either supervised or unsupervised OOD. Supervised OOD methods assume access to a labeled training dataset [9, 17, 20, 22, 28, 29, 44, 49], while unsupervised methods learn the training distribution from a set of unlabeled samples only [2, 3, 7, 19, 25, 31, 32, 38, 39, 40, 41, 47]. After training, these methods are evaluated on a test set once and the final performance is given. More closely related to the present work is the setting of OOD detection for continual classification [14]. However, our work focuses on continuous evaluation of (U-)OOD detection, which, to the best of our knowledge, has not been explored thus far. We describe related fields below, along with their differences with respect to continual U-OOD.
Test-time adaptation in OOD (TTA-OOD) is closely related to our work in that the entire unlabeled test set is used to update an OOD detector, and the performance is assessed after processing all samples. TTA-OOD works include AdaODD [10], which keeps a memory bank of features that is updated using confident pseudo-in and pseudo-OOD samples to score subsequent test samples. Another approach, ETLT [10], employs linear regression to predict the output of an OOD detector. Similarly, [8, 16, 23, 56] assume a large set of unlabeled data consisting of in- and OOD samples is available after deployment.
Crucially, TTA-ODD works only consider the setting with thousands of test samples readily available rather than investigating the progression of OOD detectors over time. As a result, they differ from our setting in two important ways. First, assuming a large test set with many OOD samples is unrealistic due to their relative scarcity. Second, evaluating OOD detectors only after a large amount of data is collected rather than its evolution ignores the more realistic and challenging early stage of exposure to OOD samples. Additionally, all these methods require access to the original labeled training set.
Outlier exposure [18] uses a large auxiliary dataset of samples that can be considered outliers in conjunction with the training data to train an improved OOD detector. OE has been successfully applied to increase the performance of many OOD detectors [11, 18, 19, 36, 38, 42, 44], but its usefulness is limited when the OE distribution is far from the OOD distribution [39]. Related to our work, [30] find that a simple classification approach between in-distribution and OE data is effective for deep anomaly detection on images and note that by increasing the amount of OE data, an unsupervised approach should transition into a supervised approach. However, they do not investigate how to make this transition, nor do they use unlabeled test samples but rather confirmed OE outliers.
OOD with limited labels. Several fields deal with OOD detection in the presence of limited labels. For instance, few-shot OOD methods assume access to a small number of labeled OOD samples [11, 34, 47]. Other works consider the case where a handful of samples per class in the training set are available[6, 12]. In semi-supervised OOD, a small set of labeled anomalies is available next to a sizeable unlabeled set of normal data [16, 35, 43, 56]. Continual U-OOD can be seen as a sequence of few-shot OOD experiments with imperfect labels obtained from the previous model.
3 Continual U-OOD
@@
The goal of U-OOD detection is to detect OOD samples and prevent them from being processed by a downstream model. In this setting, the training distribution of the downstream model is often called the in distribution $\mathcal{P}_{\textrm{in}}$, and the out distribution $\mathcal{P}_{\textrm{out}}$ is assumed to be its complement. Given its large support, producing a set of OOD samples representative of $\mathcal{P}_{\textrm{out}}$ for supervised classification of ID vs. OOD samples is intractable. Instead, U-OOD detection methods rely on in-distribution (ID) samples to train a detector $\sigma^{\textrm{in}}\colon\mathcal{X}\to$ that scores the OOD-ness of test samples at inference time. Moreover, the likelihood of observing a specific image from  $\mathcal{P}_{\textrm{out}}$ is not uniform once a downstream task is established and the model is deployed in a given environment. Instead, OOD samples detected during the operation of the deployed system, which depends on the downstream task and the deployment environment, are representative and constrain  $\mathcal{P}_{\textrm{out}}$ to the specific application. The goal of continual U-OOD is to progressively enrich the detector $\sigma^{\textrm{in}}$ with the observed OOD samples, thus adapting the detection model to the realities of the deployment environment.
@@
More specifically, a continual U-OOD system builds a sequence of detectors $s_{t}(\mathbf{x})\colon\mathcal{X}\mapsto$ for time steps $t\in\{0,1,2,\ldots\}$. The sequence starts with the base U-OOD detector, $s_{0}=\sigma^{\textrm{in}}$, and progressively adapts to the OOD samples observed after deployment. At each time $t$, the detector $s_{t}$ is trained with the dataset of samples observed until the time step $t$, denoted by $\mathcal{D}_{t}=\mathcal{D}^{\textrm{train}}\cup\mathcal{D}_{t}^{\textrm{deploy}}$. The training dataset $\mathcal{D}^{\textrm{train}}$ contains the ID samples used to train both the downstream model and the base U-OOD detector $\sigma^{\textrm{in}}$, while the deployment dataset $\mathcal{D}_{t}^{\textrm{deploy}}$ contains ID and OOD samples observed during the operation of the deployed system until time $t$. We assume that $\mathcal{D}_{0}^{\textrm{deploy}}=\varnothing$.
Training detector $s_{t}$ with $\mathcal{D}_{t}$ is a supervised process, but not all elements in $\mathcal{D}_{t}$ are labeled as ID or OOD. That is, elements of $\mathcal{D}^{\textrm{train}}$ are known to come from $\mathcal{P}_{\textrm{in}}$ and are, therefore, ID samples. In contrast, the elements of $\mathcal{D}_{t}^{\textrm{deploy}}$ are samples of the deployment distribution $\mathcal{P}_{\textrm{deploy}}$ that, following [8, 23], can be modeled with the Huber contamination model [21],
@@
with OOD sample ratio $\pi$. Consequently, the nature of the elements in $\mathcal{D}_{t}^{\textrm{deploy}}$ is unknown. We leverage the sequential nature of the detectors and use $s_{t-1}$ to pseudo-label the elements of $\mathcal{D}_{t}^{\textrm{deploy}}$ and train $s_{t}$ with them. At each time $t$, the dataset $\mathcal{D}_{t}$ can be split into two disjoint subsets $\mathcal{D}_{t}^{\textrm{in}}$ and $\mathcal{D}_{t}^{\textrm{out}}$ containing the samples labeled as ID and the samples labeled as OOD, respectively.
For simplicity and without loss of generality, we assume that the deployment dataset $\mathcal{D}_{t}^{\textrm{deploy}}$ grows $K$ elements at each time step, whereby $\left|\mathcal{D}_{t}^{\textrm{deploy}}\right|=t\cdot{}K$. We will also omit the index $t$ where not explicitly needed.
@@ 3.1 Model
The main challenge of OOD detection is the scarcity of representative OOD samples. In continual U-OOD detection, the expected number of OOD samples at time step $t$ is $\pi\cdot{}t\cdot{}K$, which for a small $t$ is too small to effectively train a binary classifier, preventing us from just using binary classifiers to model the detector $s_{t}$. As $t$ increases, however, the feasibility of training a binary classifier improves. We, therefore, design our detection model $s_{t}$ to behave as a few-shot learner $s^{-}$ when $t$ is close to 00 and to gradually transition towards a strong binary learner $s^{+}$ as $t$ increases. Formally, we model the detector $s_{t}$ as a convex combination of two learner modalities controlled by a mixing factor $\alpha_{t}$,
@@
The key difference between the learners $s^{-}$ and $s^{+}$ relies on their inductive biases. The few-shot learner $s^{-}$ is a low-variance/high-bias classifier with strong assumptions about in- and out-distributions. Its design is based on top of the U-OOD detector $\sigma^{\textrm{in}}$, as detailed below. In contrast, the strong learner $s^{+}$ is a low-bias classifier, and its architecture can be chosen according to the nature of the input space $\mathcal{X}$ (e.g., a CNN or a transformer architecture for image data), as our approach is agnostic to the internal specifics of $s^{+}$. At each step $t$, both $s^{-}_{t}$ and $s^{+}_{t}$ are trained independently with the dataset $\mathcal{D}_{t}$ pseudo-labeled with $s_{t-1}$.
The factor $\alpha_{t}\in[0,1]$ controls the transition between both models and is proportional to the number of OOD elements in $\mathcal{D}_{t}$,
@@
where the factor $\beta$ is a hyperparameter of our method.
The next sections describe our U-OOD detector $\sigma$ and how it is used to define the few-shot learner $s^{-}$.
@@ 3.1.1 U-OOD detector:
Figure 2: Application of our method on toy data. A darker background indicates more OOD. In the first stage, we train our U-OOD base detector $\sigma^{in}$. During early deployment, our method relies heavily on the few-shot learner $s^{-}$, which can increase performance with a handful of pseudo-labeled test samples. Later on, given sufficient pseudo-labeled test samples providing information about $\mathcal{P}_{\textrm{out}}$, we put more weight on the binary classifier $s^{+}$.
Our U-OOD detector combines elements from the Mahalanobis anomaly detector (MahaAD) [41], known for its robustness and speed [7], and from non-parametric nearest-neighbor scoring methods [2, 38, 49, 50]. As in MahaAD, given the collection of ID samples $\mathcal{D}^{\textrm{in}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, we fit a Gaussian parameterized by the data mean $\bm{\mu}$ and data covariance matrix $\bm{\Sigma}$. To prevent numerical problems with near-singular covariance matrices in high-dimensional or low-data regimes, shrinkage is applied, resulting in the regularized covariance matrix,
$\mathbf{\Sigma}^{\prime}=(1-\nu)\mathbf{\Sigma}+\nu\frac{\operatorname{Tr}(\mathbf{\Sigma})}{d}\mathbf{I}_{d},$ (4)
where the shrinkage factor $\nu$ is found using the standard hyperparameter-free method of [27] and $d$ is the data dimensionality. MahaAD uses the induced Mahalanobis distance $d_{\mathbf{\Sigma}^{\prime}}(\mathbf{x},\bm{\mu})$ between the data mean $\bm{\mu}$ and a test sample $\mathbf{x}$ to estimate its OOD score. We instead use this distance to perform a k-NN search and score test samples with the average distance to their $k$ nearest neighbors,
$\sigma^{\textrm{in}}(\mathbf{x})=\dfrac{1}{k}\sum_{\mathbf{x}^{\prime}\in N^{k}_{\mathbf{\Sigma}^{\prime}}(\mathbf{x})}d_{\mathbf{\Sigma}^{\prime}}(\mathbf{x},\mathbf{x}^{\prime}),$ (5)
where $N^{k}_{\mathbf{\Sigma}^{\prime}}(\mathbf{x})$ denotes the k-nearest neighbors of $\mathbf{x}$ in the training data measured with the Mahalanobis distance induced by $\mathbf{\Sigma}^{\prime}$.

### v2

# Iterative Deployment Exposure for Unsupervised Out-of-Distribution Detection

@@ Abstract
Deep learning models are vulnerable to performance degradation when encountering out-of-distribution (OOD) images, potentially leading to misdiagnoses and compromised patient care. These shortcomings have led to great interest in the field of OOD detection. Existing unsupervised OOD (U-OOD) detection methods typically assume that OOD samples originate from an unconcentrated distribution complementary to the training distribution, neglecting the reality that deployed models passively accumulate task-specific OOD samples over time. To better reflect this real-world scenario, we introduce Iterative Deployment Exposure (IDE), a novel and more realistic setting for U-OOD detection. We propose CSO, a method for IDE that starts from a U-OOD detector that is agnostic to the OOD distribution and slowly refines it during deployment using observed unlabeled data. CSO uses a new U-OOD scoring function that combines the Mahalanobis distance with a nearest-neighbor approach, along with a novel confidence-scaled few-shot OOD detector to effectively learn from limited OOD examples. We validate our approach on a dedicated benchmark, showing that our method greatly improves upon strong baselines on three medical imaging modalities.
@@
Out-of-Distribution Detection Deployment Reliability.
@@ 1 Introduction
Deep learning (DL) models fundamentally rely on the premise that the training data distribution aligns with that of the test data. However, this assumption often fails in real-world situations where the performance of a DL model diverges from its initial benchmark due to encounters with out-of-distribution (OOD) samples. This phenomenon is highly problematic in medical imaging, where the dependability of DL models is critical for safe, prolonged use in the field.
Figure 1: Iterative deployment exposure. The initial detector, trained on unlabeled ID data, is iteratively refined with unlabeled deployment samples.
These shortcomings have sparked great interest in the field of OOD detection [7, 27, 29], and unsupervised OOD (U-OOD) detection in particular. Unlike supervised OOD detection, U-OOD assumes neither access to training labels nor OOD samples, thereby encompassing a more generally applicable albeit more challenging setting [2, 5, 6, 12, 14, 15]. The core principle of U-OOD is to identify the level sets of the in-distribution (ID) training data and establish a threshold to distinguish OOD samples, where the assumption is that OOD samples are the complement of the ID and therefore unconcentrated [21, 23].
However, this assumption fails to consider the underlying motivation of OOD detection, that is, improving the reliability of deployed downstream models. Treating OOD data as arising from a generic, unconcentrated distribution is a disconnect from the reality that deployment environments inherently impose a specific, concentrated OOD context. We argue that the OOD context must be inferred from the deployment setting, and, in turn, the deployment context fully determines a concentrated OOD distribution. As this context depends on the deployment, U-OOD works that synthesize anomalies a priori (e.g., [20, 25]) to simulate the OOD distribution are inadequate. Instead, an approach where an initial U-OOD detector, agnostic to the OOD distribution, is gradually updated during deployment to consider the actual OOD distribution is needed (Fig. 1).
@@
Yet, current U-OOD research has mainly neglected this important consideration. While related fields like OOD detection with in-the-wild data [3, 9] and OOD test-time adaptation [4, 28] exist, they are predicated on the availability of a large number of test samples: the former relies on a large set of unlabeled data consisting of both ID and OOD samples, while the latter uses the entire unlabeled test set to update a model. In both cases, hundreds of OOD samples are used to update their respective models, and performance is only measured after seeing all of them. Consequently, this is often unrealistic, considering that OOD samples can be scarce in medical applications.
@@
Instead, we focus on improving detection with only a few OOD samples and tackle the realistic setting where this process is iterative: detect OOD samples, update the OOD detector, and repeat the procedure with a refined model several times. Consequently, evaluating U-OOD detectors should extend beyond a single instance and consider their effectiveness over time. We refer to this setting as IDE (Iterative Deployment Exposure). Hence, in this work, we,
@@
    introduce the problem of IDE for U-OOD detection, which closely matches the reality of model deployment.
@@
    introduce new metrics and benchmarks to evaluate methods in the IDE setting.
@@
    propose CSO, a novel method for IDE that outperforms strong baselines from related fields.
2 Iterative Deployment Exposure
@@
In U-OOD detection, the training distribution of the downstream model is called the in distribution $\mathcal{P}_{\textrm{in}}$, and the out distribution $\mathcal{P}_{\textrm{out}}$ is assumed to be its complement. Given its large support, producing a set of OOD samples representative of $\mathcal{P}_{\textrm{out}}$ for supervised classification of ID vs. OOD samples is intractable. Instead, U-OOD detection methods rely on in-distribution (ID) samples to train a detector $\sigma^{\textrm{in}}\colon\mathcal{X}\to\mathbb{R}$ that scores the OOD-ness of test samples at inference time. Critically, the likelihood of observing a specific image from  $\mathcal{P}_{\textrm{out}}$ is not uniform once a downstream task is established and the model is deployed in a given environment. Instead, OOD samples seen during the operation of the deployed system constrain  $\mathcal{P}_{\textrm{out}}$ to the specific application. The goal of IDE is to progressively enrich the detector $\sigma^{\textrm{in}}$ with observed OOD samples, thus adapting the detection model to the deployment environment.
@@
More specifically, an IDE system builds a sequence of detectors $s_{t}(\mathbf{x})\colon\mathcal{X}\mapsto\mathbb{R}$ for time steps $t\in\{0,1,2,\ldots\}$. The sequence starts with the base U-OOD detector, $s_{0}=\sigma^{\textrm{in}}$, and progressively adapts to the OOD samples observed after deployment. At each time $t$, the detector $s_{t}$ is trained with the dataset of samples observed until time step $t$, denoted $\mathcal{D}_{t}=\mathcal{D}^{\textrm{train}}\cup\mathcal{D}_{t}^{\textrm{deploy}}$. We assume that $\mathcal{D}_{0}^{\textrm{deploy}}=\varnothing$. Crucially, not all elements in $\mathcal{D}_{t}$ are labeled as ID or OOD. While elements of $\mathcal{D}^{\textrm{train}}$ are known to come from distribution $\mathcal{P}_{\textrm{in}}$ and are, therefore, ID samples, elements of $\mathcal{D}_{t}^{\textrm{deploy}}$ are unlabeled. These samples of the deployment distribution $\mathcal{P}_{\textrm{deploy}}$, following [3, 9], are modeled with the Huber contamination model [8],
@@
with contamination ratio $\pi$.
To address the lack of labels in $\mathcal{D}_{t}^{\textrm{deploy}}$$s_{t-1}$ is used to pseudo-label the elements of $\mathcal{D}_{t}^{\textrm{deploy}}$ for training $s_{t}$. Hence, at each time $t$$\mathcal{D}_{t}$ can be split into two disjoint subsets $\mathcal{D}_{t}^{\textrm{in}}$ and $\mathcal{D}_{t}^{\textrm{out}}$ containing the samples labeled as ID and OOD, respectively. For simplicity and without loss of generality, we assume that the deployment dataset $\mathcal{D}_{t}^{\textrm{deploy}}$ grows $K$ elements at each time step, whereby $\left|\mathcal{D}_{t}^{\textrm{deploy}}\right|=t\cdot{}K$. We will also omit the index $t$ where not explicitly needed.
@@ 2.1 Model
The main challenge of OOD detection is the scarcity of representative OOD samples. In IDE, the expected number of OOD samples at time step $t$ is $\pi\cdot{}t\cdot{}K$, which for a small $t$ is too small to effectively train a binary classifier, preventing us from simply using binary classifiers to model the detector $s_{t}$. As $t$ increases, however, the feasibility of training a binary classifier improves. We, therefore, design our detection model $s_{t}$ to behave as a few-shot learner $s^{-}$ when $t$ is close to 00 and to gradually transition towards a strong binary learner $s^{+}$ as $t$ increases. Formally, we model the detector $s_{t}$ as a convex combination of two learner modalities controlled by a mixing factor $\alpha_{t}$,
@@
The key difference between the learners $s^{-}$ and $s^{+}$ lies in their inductive biases. The few-shot learner $s^{-}$ is a low-variance/high-bias classifier with strong assumptions about in- and out-distributions. Its design is based on the U-OOD detector $\sigma^{\textrm{in}}$, as detailed below. In contrast, the strong learner $s^{+}$ is a low-bias binary classifier, and its architecture can be chosen according to the nature of the input space $\mathcal{X}$ (e.g., a CNN or a transformer architecture for image data), as our approach is agnostic to the internal specifics of $s^{+}$. At each step $t$, both $s^{-}_{t}$ and $s^{+}_{t}$ are trained independently with the dataset $\mathcal{D}_{t}$ pseudo-labeled with $s_{t-1}$.
The factor $\alpha_{t}\in[0,1]$ controls the transition between both models and is proportional to the number of elements pseudo-labeled as OOD,
@@
where the factor $\beta$ is a hyperparameter of our method. We refer to our method as CSO (confidence-scaled U-OOD detector). The next sections describe our U-OOD detector $\sigma^{in}$ and how it is used to define the few-shot learner $s^{-}$.
@@ 2.1.1 U-OOD detector:
Our U-OOD detector combines elements from the Mahalanobis anomaly detector (MahaAD) [19], known for its robustness and speed [2], and from non-parametric nearest-neighbor scoring methods [1, 17, 24]. As in MahaAD, given the collection of ID samples $\mathcal{D}^{\textrm{in}}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, we fit a Gaussian distribution parameterized by the data mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$. To prevent numerical problems with near-singular covariance matrices in high-dimensional or low-data regimes, we use shrinkage following the standard hyperparameter-free method of [11]. MahaAD uses the induced Mahalanobis distance $d_{\mathbf{\Sigma}^{\prime}}(\mathbf{x},\bm{\mu})$ between $\bm{\mu}$ and a test sample $\mathbf{x}$ to estimate its OOD score. We instead use this distance to perform a 2-NN search (following, e.g., [1]) and score test samples with the average distance to their $2$ nearest neighbors,
$\sigma^{\textrm{in}}(\mathbf{x})=\dfrac{1}{2}\sum_{\mathbf{x}^{\prime}\in N^{2}_{\mathbf{\Sigma}^{\prime}}(\mathbf{x})}d_{\mathbf{\Sigma}^{\prime}}(\mathbf{x},\mathbf{x}^{\prime}),$ (4)
where $N^{2}_{\mathbf{\Sigma}^{\prime}}(\mathbf{x})$ denotes the 2-nearest neighbors of $\mathbf{x}$ in the training data measured with the Mahalanobis distance induced by $\mathbf{\Sigma}^{\prime}$. We refer to this as MkNN.